# Privy: Envisioning and Mitigating Privacy Risks for Consumer-facing AI Product Concepts

### v1

@@ Abstract.
AI creates and exacerbates privacy risks, yet practitioners lack effective resources to identify and mitigate these risks. We present Privy, a tool that guides practitioners through structured privacy impact assessments to: (i) identify relevant risks in novel AI product concepts, and (ii) propose appropriate mitigations. Privy was shaped by a formative study with 11 practitioners, which informed two versions — one LLM-powered, the other template-based. We evaluated these two versions of Privy through a between-subjects, controlled study with 24 separate practitioners, whose assessments were reviewed by 13 independent privacy experts. Results show that Privy helps practitioners produce privacy assessments that experts deemed high quality: practitioners identified relevant risks and proposed appropriate mitigation strategies. These effects were augmented in the LLM-powered version. Practitioners themselves rated Privy as being useful and usable, and their feedback illustrates how it helps overcome long-standing awareness, motivation, and ability barriers in privacy work.
@@ 1. Introduction
Practitioners building AI products often lack awareness of how privacy risks manifest, are not incentivized to address them, and have few resources to guide effective mitigation (Lee et al., 2024a). While prior work has proposed taxonomies to characterize emerging privacy threats in AI systems (Lee et al., 2024b; Solove, 2025; Windl et al., 2023), and broad guidelines for building responsible AI products that address some privacy concerns (Amershi et al., 2019; PAIR, 2019; Apple, 2023), there remains a critical tooling gap. Practitioners lack support for applying these frameworks to early-stage product development. This reflects the broader “principle-practice gap” in ethical and responsible AI — guidelines are plentiful but rarely translate into actionable design processes (Winfield and Jirotka, 2018; Shneiderman, 2020). In this paper, we ask: How can we design tools that make it easier for practitioners to identify and mitigate privacy risks in their AI products and services?
@@
Canonically, privacy experts conduct “privacy impact assessments” (PIAs) to identify privacy risks with novel product concepts (Clarke, 2009; Wong and Mulligan, 2019). PIAs take the form of a structured report that prompts experts to think about how the design of a product may cause privacy risks, and how those risks might be mitigated (noa, 2023; van Puijenbroek and Hoepman, 2017). However, preparing these reports requires strong subject matter expertise in privacy and is labor-intensive, preventing practitioners who are not privacy experts from participating in the development of these reports. Moreover, PIAs are slow to evolve and many have not been updated to account for the new and emerging risks of AI technologies.
@@
More recently, researchers have explored the use of Large Language Model (LLM) technologies to help practitioners surface ethical concerns and harms (Wang et al., 2024; Buçinca et al., 2023; Herdel et al., 2024). Yet, no work has produced artifacts tailored to privacy risks or examined how LLM-powered systems might support practitioners in both identifying and mitigating them. Building off this prior work, we developed Privy, a privacy risk-envisioning tool that helps AI practitioners — regardless of their privacy expertise — identify and mitigate AI privacy risks in novel product concepts.
@@
We first adapted Farsight (Wang et al., 2024), a popular LLM-powered risk envisioning system, to the privacy domain using the AI privacy taxonomy (Lee et al., 2024b). With this early prototype, we conducted a formative study with 11 AI and privacy practitioners to distill key design goals. We then implemented two versions of Privy: 1) Privy-Template: a non-AI, static template-based version, and 2) Privy-LLM: an LLM-powered version that offers interactive suggestions. Both versions follow the same risk-envisioning workflow. We evaluated Privy with 24 AI practitioners (between-subjects: LLM vs. Template), whose resulting PIA reports were assessed by 13 privacy experts.
@@
    How and to what extent do Privy-LLM and Privy-Template help practitioners identify and mitigate privacy risks with AI product concepts?
@@
Our findings show that Privy effectively scaffolds practitioners in creating high-quality PIAs. Using both versions of Privy, practitioners identified risks that were both relevant and severe, and developed effective and appropriate mitigation strategies (RQ1, Section 6.1). Compared to Privy-Template, Privy-LLM produced assessments in which the identified risks were judged by experts to be more relevant, severe, and correct, with clearer articulation. Privy-LLM also produced mitigation plans that were more effective and appropriate for the identified risks. Yet both versions were seen as comparatively useful and usable (RQ2, Section 6.2). Privy also overcame long-standing barriers that hinder practitioners’ privacy work by improving awareness (via scaffolding risk identification), fostering motivation (via encouraging reflection and engagement), and supporting ability (via enabling self-efficacy privacy practice) (RQ3, Section 6.3).
@@
  • The design and implementation of Privy, both Privy-LLM and Privy-Template, that instantiate these design goals (Section 4).
  • Empirical insights from an evaluation study with 24 AI practitioners and 13 privacy experts, demonstrating that Privy effectively scaffolds creating quality PIAs (Section 6.1), that LLMs enhance assessment quality through human–AI collaboration (Section 6.2), and that Privy helps practitioners overcome awareness, motivation, and ability barriers (Sections 6.3).
@@ 2.2. Practitioners’ approaches to and challenges with addressing privacy risks
A growing body of research investigates how AI practitioners grapple with privacy risks during the development and deployment of AI systems. Although many practitioners recognize privacy as a critical concern, they often struggle to address privacy risks due to organizational constraints, a lack of supportive policies, and insufficient tools (Ali et al., 2023; Lee et al., 2024a). Pant et al. (Pant et al., 2023) identify notable challenges in developing ethical and privacy-compliant AI systems: limited training, fragmented understanding of privacy, inadequate foresight, and organizational inertia.
One proposed solution is adopting a privacy-by-design approach, which parallels human-centered design in advocating cross-functional integration of privacy considerations throughout the development lifecycle (Al Breiki and Mahmoud, 2025). However, in practice, privacy work is often siloed, delegated to legal or compliance teams, rather than integrated into product workflows (Lee et al., 2024a; Meisenbacher et al., 2024). This handoff model can be inefficient, leading to additional rework cycles as teams revise products to meet regulatory expectations.
@@
These realities underscore the need for tools that support AI development to ensure that privacy becomes a fundamental element rather than an afterthought. Answering this call, we present Privy (Section 4), the first artifact specifically designed to help practitioners, who are not privacy experts, foreground privacy assessment for early-phase AI product concepts. Section 3 details our formative study with AI and privacy practitioners to understand their needs for such a tool. Section 5 describes how we tested Privy with real-world AI practitioners to evaluate its impact on their privacy approach in AI product development.
2.3. Tooling for AI risk assessment
@@
Various tools have emerged to help practitioners identify and assess the risks posed by AI systems. In general, these fall into two categories: structured worksheets and interactive interfaces.
@@
Structured worksheets are static documents that prompt reflection on potential harms. These include privacy checklists that guide developers through risk considerations across the AI lifecycle (Deng et al., 2025; Bogucka et al., 2024b), and consequence scanning exercises that help teams anticipate both the intended and unintended impact of AI on users (doteveryone, 2019). Government frameworks, such as NIST’s AI RMF (of Standards and Technology, 2023), similarly promote responsible documentation of development decisions and mitigations.
One common tool is the Privacy Impact Assessment (PIA) (Clarke, 2009), which helps record design decisions and surface privacy risks. However, PIAs often fall short in connecting identified risks to the AI system’s capabilities, requirements, and use cases. They rarely provide a comprehensive view of the system’s information flow, for example, how one use case can generate both intended and unintended consequences (Lee et al., 2024a).
@@
Interactive interfaces build upon the principles of structured worksheets but offer digital, dynamic workflows. Some tools integrate stakeholder input with internal privacy expertise to facilitate impact assessments (Bogucka et al., 2024a), while others use card-based prompts to encourage teams to consider potential effects on diverse stakeholders (Elsayed-Ali et al., 2023). These tools improve awareness, but still face key limitations: they are slow to adapt, rarely cover the full AI lifecycle, and often lack flexibility to evolving regulatory or technical contexts (Kuehnert et al., 2025).
@@
To address these gaps, Privy is designed to guide practitioners through use case ideation, privacy risk identification, and risk mitigation planning, while clearly linking risks to AI system capabilities. Drawing from both structured assessments and interactive tools, we implemented two versions: 1) Privy-Template, a structured worksheet administered as a Google Doc template, and 2) Privy-LLM, an interactive interface enhanced with LLM-generated suggestions.
@@ 2.4. Impacts of GenAI-powered tooling on AI practitioners’ privacy work
3. Formative Study and Design Goals
We developed an early prototype, Privy beta, by adapting Farsight (Wang et al., 2024), a general AI risk-envisioning system, to the context of AI privacy risks (Lee et al., 2024b). Privy beta included two core interfaces (Appendix C.1): (1) the risk explorer, which guided users through articulating a product’s purpose, use cases, and impacted stakeholders to identify the most pertinent of 12 AI privacy risks; and (2) the risk summarizer, which synthesized users’ identified risks and surfaced related real-world incidents sourced from the AIAAIC, a popular AI incident database (Pownall, 2023).
Using Privy beta, we conducted a formative study with $N=11$ practitioners. We asked: how can we design Privy to better support practitioners in identifying AI privacy risks and making privacy-informed design decisions for novel product concepts? From this study, we derived five design goals (DGs) for Privy (DG1–DG5; Section 4).
@@ 3.1. Participants
We recruited 11 AI and privacy practitioners through direct contacts and social networks, spanning eight companies (nine in the US, two in the UK). All had experience designing or developing consumer-facing AI products. Each 60-minute session was compensated with a \$100 Amazon gift card. Participants included eight men and three women with diverse roles (Appendix C.2). Nine reported having engaged in end-user privacy discussions as part of their work.
@@ 3.2. Procedure
Each session began with participants describing their experience identifying privacy risks for an early-stage consumer-facing AI product they had worked on. After a walkthrough of Privy beta, they used the risk explorer to identify risks in that AI product and then reviewed a one-page privacy report in the risk summarizer. Participants were asked to think aloud during the session and provide feedback on Privy beta’s features. We concluded with a semi-structured interview (Appendix C.3) where participants reflected on Privy beta’s potential fit in their work and suggested alternative design ideas.
All study sessions were audio- and video-recorded and then transcribed. Two researchers coded the transcripts independently, and later came together to resolve any conflicts in coding and identify emerging use patterns and themes.
@@ 3.3.1. Scaffold product-specific AI privacy implications
Overall, practitioners benefited from a structured workflow that links AI product concepts to privacy risks and valued tailored support to guide them through it.

### v2

@@ Abstract.
AI creates and exacerbates privacy risks, yet practitioners lack effective resources to identify and mitigate these risks. We present Privy, a tool that guides practitioners without privacy expertise through structured privacy impact assessments to: (i) identify relevant risks in novel AI product concepts, and (ii) propose appropriate mitigations. Privy was shaped by a formative study with 11 practitioners, which informed two versions — one LLM-powered, the other template-based. We evaluated these two versions of Privy through a between-subjects, controlled study with 24 separate practitioners, whose assessments were reviewed by 13 independent privacy experts. Results show that Privy helps practitioners produce privacy assessments that experts deemed high quality: practitioners identified relevant risks and proposed appropriate mitigation strategies. These effects were augmented in the LLM-powered version. Practitioners themselves rated Privy as being useful and usable, and their feedback illustrates how it helps overcome long-standing awareness, motivation, and ability barriers in privacy work.
Privacy, Human-centered AI, Privacy Impact Assessment, Human-AI Collaboration
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791279isbn: 979-8-4007-2278-3/2026/04ccs: Security and privacy Human and societal aspects of security and privacyccs: Human-centered computing Human computer interaction (HCI)
@@
Screenshot of the Privy-LLM interface, which supports AI practitioners in identifying and mitigating privacy risks through a structured, node-based workflow. (A) The Product Description Node (leftmost, green) allows users to describe their AI product concept. (B-1) The Use Case Node (blue) lets users input intended and unintended use cases, while (B-2) the Use Case Suggestions panel displays additional AI-generated examples when users click “GET SUGGESTIONS.” (C) The AI Capabilities and Requirements Node (green) summarizes the AI’s inferred capabilities and system needs based on described use cases. (D-1) The Privacy Risk Assessment Node (orange) guides users to articulate how a privacy risk may arise, identify impacted stakeholders, and rate the risk’s relevancy and severity. (D-2) Risk Suggestions are shown when users request AI-generated privacy risks by clicking “GET SUGGESTIONS.” (E) The Risk Definitions Panel (rightmost, gray) provides categorized definitions of privacy risks — grouped under data collection, processing, dissemination, and physical invasion — with icons, descriptions, and external links to real-world examples. Arrows and dotted lines illustrate how nodes are interconnected, mapping the progression from product concept to privacy risk identification.
@@ 1. Introduction
Practitioners building AI products often lack awareness of how privacy risks manifest, are not incentivized to address them, and have few resources to guide effective mitigation (Lee et al., 2024a). Although prior work has proposed taxonomies to characterize emerging privacy threats in AI systems (Lee et al., 2024b; Solove, 2025; Windl et al., 2023), and broad guidelines for building responsible AI products that address some privacy concerns (Amershi et al., 2019; PAIR, 2019; Apple, 2023), there remains a critical tooling gap. Practitioners often lack both privacy expertise and the support needed to apply these frameworks to early stage product development. This reflects the broader “principle-practice gap” in ethical and responsible AI — guidelines are plentiful but rarely translate into actionable design processes (Winfield and Jirotka, 2018; Shneiderman, 2020). In this paper, we ask: How can we design tools that make it easier for practitioners to identify and mitigate privacy risks in their AI products and services?
@@
Canonically, privacy experts conduct “privacy impact assessments” (PIAs) to identify privacy risks with novel product concepts (Clarke, 2009; Wong and Mulligan, 2019). PIAs take the form of a structured report that prompts experts to think about how the design of a product may cause privacy risks, and how those risks might be mitigated (34; J. van Puijenbroek and J. Hoepman (2017)). However, existing PIA frameworks (e.g., ISO’s Information Technology Security Techniques (34)) have not been substantially updated to account for the evolving and emerging privacy risks introduced by AI technologies (Lee et al., 2024b). Moreover, preparing these reports requires strong privacy expertise and is labor-intensive, precluding participation by contributing practitioners who are not themselves privacy experts.
@@
More recently, researchers have explored the use of Large Language Model (LLM) technologies to help practitioners surface ethical concerns and harms (Wang et al., 2024; Buçinca et al., 2023; Herdel et al., 2024). However, no work has produced LLM-powered systems that might support practitioners in both identifying and mitigating privacy risks. Building off this prior work, we developed Privy, a privacy risk-envisioning tool that helps AI practitioners — even those with little privacy expertise — identify and mitigate AI privacy risks in novel product concepts.
@@
To inform the design of Privy, we first implemented an early prototype by adapting Farsight, a general-purpose AI risk envisioning workflow (Wang et al., 2024) from the responsible AI literature, to the privacy domain using the AI privacy risk taxonomy (Lee et al., 2024b). With this early prototype, we conducted a formative study with 11 AI and privacy practitioners to distill key design goals. In short, Privy guides AI practitioners through a structured workflow to: (i) elicit the capabilities and requirements of their AI product concepts (Fig. 1A-C), (ii) identify product-specific privacy risks (Fig. 1D), and (iii) brainstorm pertinent mitigation strategies (Fig. 3). We implemented two versions of Privy: 1) PrivyTemplate: a non-AI, static template-based version, and 2) PrivyLLM: an LLM-powered version that offers suggestions interactively. Both versions follow the same risk-envisioning workflow. We evaluated Privy with 24 non-privacy-expert AI practitioners (between-subjects: LLM vs. Template), whose resulting PIA reports were assessed by 13 privacy experts.
@@
    How and to what extent do PrivyLLM and PrivyTemplate help practitioners identify and mitigate privacy risks with AI product concepts?
@@
Our findings show that Privy effectively scaffolds non-privacy-expert practitioners in creating high-quality PIAs. Using both versions of Privy, practitioners identified risks that were relevant and severe, and developed effective and appropriate mitigation strategies (RQ1, Section 6.1). Compared to PrivyTemplate, PrivyLLM produced assessments in which the identified risks were judged by privacy experts to be more relevant, severe, and correct, with clearer articulation. PrivyLLM also produced mitigation plans that were more effective and appropriate for the identified risks. Both versions were seen as useful and usable (RQ2, Section 6.2). Privy also overcame long-standing barriers that hinder non-privacy-expert practitioners’ privacy work by improving awareness (via scaffolding risk identification), fostering motivation (via encouraging reflection and engagement), and supporting ability (by improving privacy self-efficacy) (RQ3, Section 6.3).
@@
  • The design and implementation of Privy, both PrivyLLM and PrivyTemplate, that instantiate these design goals (Section 4).
  • Empirical insights from an evaluation study with 24 non-privacy-expert AI practitioners and 13 privacy experts, demonstrating that Privy effectively scaffolds creating quality PIAs (Section 6.1), that LLMs enhance assessment quality through human–AI collaboration (Section 6.2), and that Privy helps practitioners overcome awareness, motivation, and ability barriers in privacy work (Sections 6.3).
@@ 2.2. Practitioners’ approaches to and challenges with addressing privacy risks
A growing body of research investigates how AI practitioners grapple with privacy risks during the development and deployment of AI systems. Although many practitioners recognize privacy as a critical concern, they often struggle to address privacy risks due to organizational constraints, a lack of supportive policies, and insufficient tools (Ali et al., 2023; Lee et al., 2024a; Das et al., 2023). Pant et al. (Pant et al., 2023) identify notable challenges in developing ethical and privacy-compliant AI systems: limited training, fragmented understanding of privacy, inadequate foresight, and organizational inertia.
One line of work extends practices from requirements engineering (Herwanto et al., 2024; Gharib et al., 2016) to operationalize privacy by design (Al Breiki and Mahmoud, 2025; Gürses et al., 2011) principles, giving rise to what is now termed “privacy engineering”. Within software engineering, this approach conceptualizes two levels at which privacy can be incorporated into a system: privacy by policy and privacy by architecture (Spiekermann and Cranor, 2009). These levels reflect different mechanisms for embedding privacy in product design workflows, be it through user-facing notice and consent controls or through constraining data collection and processing at the product-level.
Yet in practice, privacy work and responsibility are often organizationally siloed. Engineering teams are tasked with executing decisions handed down from legal or compliance teams (Gürses and del Alamo, 2016; Kostova et al., 2020), rather than collaboratively working with these teams to define what privacy should mean for a product and why (Lee et al., 2024a; Meisenbacher et al., 2024). This handoff model can introduce inefficiencies and additional rework as teams iterate to align products with regulatory expectations.
@@
These realities underscore the need for tools that support AI development to ensure that privacy becomes a fundamental element rather than an afterthought. Answering this call, we present Privy (Section 4), the first artifact specifically designed to help practitioners, who are not privacy experts, foreground privacy impact assessment for early-phase AI product concepts. Section 3 details our formative study with AI and privacy practitioners to understand their needs for such a tool. Section 5 describes how we tested Privy with real-world AI practitioners to evaluate its impact on their privacy approach in AI product development.
2.3. Tooling for privacy risk assessment
@@
Various tools have emerged to help practitioners identify and assess privacy risks. In general, these fall into two categories: structured worksheets and interactive interfaces.
Structured worksheets are static documents that prompt reflection on potential harms. The dominant tool is the Privacy Impact Assessment (PIA) (Clarke, 2009), which helps record design decisions and surface privacy risks. These assessments are intentionally technology-agnostic and originate from the need to ensure systematic documentation aligned with legal and regulatory expectations (Wright, 2013). However, PIAs are frequently regarded as difficult to operationalize, as their guidance on identifying and mitigating privacy risks is often generic and insufficiently tailored to specific technical contexts (Ahmadian et al., 2018; Ferra et al., 2019). These limitations become even more pronounced when PIAs are applied to AI-driven systems, where complexity and opacity challenge the effectiveness of traditional assessment approaches (Gilbert et al., 2025). They rarely provide a comprehensive view of how product data flows might entail both intended and unintended consequences (Lee et al., 2024a). Nonetheless, ongoing legal and privacy scholarship highlights opportunities to adapt PIAs to be more system-specific and context-aware, thereby improving their relevance and utility in AI-focused environments (Wairimu et al., 2024; Poullet, 2023). PIA worksheets directly informed the privacy-focused approach to Privy, but we contribute to the literature by meeting the calls for AI-specific privacy risks tied to the technology’s capabilities and requirements.
@@
Recently developed variations on the worksheets include privacy checklists that guide developers through risk considerations throughout the AI lifecycle (Deng et al., 2025; Bogucka et al., 2024b), and consequence scanning exercises that help teams anticipate both the intended and unintended impacts of AI on users (doteveryone, 2019). International standards bodies have also developed frameworks, such as NIST’s AI RMF (of Standards and Technology, 2023), the ISO’s Information Technology Security Techniques (29134) (34) and the Artificial Intelligence Set (42001) (35), to promote responsible documentation of development decisions and mitigations. Yet these standards are often static and driven by policy requirements rather than practitioner needs.
@@
Interactive interfaces build upon the principles of structured worksheets, but offer digital, dynamic workflows. Some tools integrate stakeholder input with internal privacy expertise to facilitate impact assessments (Bogucka et al., 2024a), while others use card-based prompts to encourage teams to consider potential effects on diverse stakeholders (Elsayed-Ali et al., 2023). Card-based approaches built on the LINDDUN methodology (Deng et al., 2011) to privacy threat analysis have been shown to lower the barrier of entry for non-privacy-expert practitioners (Wuyts et al., 2020). LINDDUN also offers other methodologies to conduct PIAs for both expert and non-expert privacy practitioners [^1] https://linddun.org/. These tools improve awareness, but still face key limitations: they are slow to adapt, rarely cover the full AI lifecycle (or AI at all), and often lack flexibility in accounting for evolving regulatory or technical contexts (Kuehnert et al., 2025).
@@
To address these gaps, Privy is designed to guide practitioners through use case ideation, privacy risk identification, and risk mitigation planning, while clearly linking risks to AI system capabilities. Drawing from both structured assessments and interactive tools, we implemented two versions: 1) PrivyTemplate, a structured worksheet administered as a Google Doc template, and 2) PrivyLLM, an interactive interface enhanced with LLM-generated suggestions.
@@ 2.4. Impacts of GenAI-powered tooling on AI practitioners’ privacy work
3. Formative Study
Since there is no established standard for AI privacy impact assessments (Section 2.3), we explored the design opportunities for this new tool through formative design engagements with an early prototype, Privybeta. We first created an interface for Privybeta that adapted the “use cases $\rightarrow$ stakeholders $\rightarrow$ harms workflow from Farsight (Wang et al., 2024) — a general, LLM-powered AI risk-envisioning system — to the context of AI privacy impact assessments. Specifically, Privybeta included two core LLM-powered interactive interfaces for practitioners to generate AI privacy impact assessments for AI product concepts: (1) the risk explorer, which guided users through articulating a product concept’s capability and requirement, use cases, and impacted stakeholders to identify the most pertinent of 12 types of AI privacy risks modeled from the privacy taxonomy (Lee et al., 2024b); and (2) the risk summarizer, which synthesized users’ identified risks and surfaced related real-world incidents sourced from the AIAAIC, a popular AI incident database (Pownall, 2023). The two interfaces can be found in Appendix C.1.
We used this early prototype to inform Privy’s design by conducting a formative study with $N=11$ practitioners. We asked: how can we design Privy to better support practitioners in identifying AI privacy risks and making privacy-informed design decisions for novel product concepts? From this study, we derived five design goals (DGs) that guided the final design of Privy (DG1–DG5; Section 4).
@@ 3.1. Participants
We recruited 11 AI and privacy practitioners (nine in the US, two in the UK) spanning eight companies through the authors’ direct contacts and social networks. We required participants to have experience with designing and/or developing consumer-facing AI products at work. Each 60-minute session was compensated with a \$100 Amazon gift card. Participants included eight men and three women with diverse roles (Appendix C.2). Nine reported having engaged in end-user privacy discussions as part of their work.
@@ 3.2. Procedure
Each session began with participants describing their experience identifying privacy risks for an early-stage consumer-facing AI product they had worked on. After being introduced to Privybeta, they used the risk explorer to identify risks in that AI product and then reviewed a one-page privacy report in the risk summarizer. Participants were asked to think aloud during the session and provide feedback on Privybeta’s features. We concluded with a semi-structured interview (Appendix C.3) where participants reflected on Privybeta’s potential fit in their product development and suggested alternative design ideas.
All study sessions were audio- and video-recorded and then transcribed. Two researchers applied iterative open coding (Service, 2009), in which they coded the transcripts independently, and met regularly to resolve any conflicts in coding and identify emerging use patterns and themes.
@@ 3.3.1. Scaffold product-specific AI privacy implications
Overall, practitioners benefited from a structured workflow that links AI product concepts to privacy risks and valued structured support to guide them through it.